# VideoPASTA [Uncaptioned image]: 7K Preference Pairs That Matter for Video-LLM Alignment

### v1

@@ Abstract
Video-language models (Video-LLMs) excel at understanding video content but struggle with spatial relationships, temporal ordering, and cross-frame continuity. To address these limitations, we introduce VideoPASTA  [Uncaptioned image] (Preference Alignment with Spatio-Temporal-Cross Frame Adversaries), a framework that enhances Video-LLMs through targeted preference optimization. VideoPASTA trains models to distinguish accurate video representations from carefully generated adversarial examples that deliberately violate spatial, temporal, or cross-frame relations. By applying Direct Preference Optimization to just 7,020 preference pairs, VideoPASTA learns robust representations that capture fine-grained spatial relationships and long-range temporal dynamics. Experiments on standard video benchmarks show significant relative performance gains of 3.05% on VideoMME, 1.97% on NeXTQA, and 1.31% on LongVideoBench, over the baseline Qwen2.5-VL model. These results demonstrate that targeted alignment, rather than massive pretraining or architectural modifications, effectively addresses core video-language challenges. Notably, VideoPASTA achieves these improvements without human annotation or captioning, relying on just 32-frame sampling, compared to the 96-frame, multi-GPU setups of prior work. This efficiency makes our approach a scalable, plug-and-play solution that seamlessly integrates with existing models while preserving their capabilities.
@@ 1 Introduction
Recent advances in video large language models (Video-LLMs) have enabled efficient understanding and reasoning about videos [65, 30, 26, 54, 55, 3, 7]. These models capture both fine-grained details and broader contextual information in videos, achieving high performance in tasks such as captioning and question answering [41, 6, 32]. However, these models typically require large, high-quality annotated datasets and substantial computational resources for training. Recent work has explored instruction tuning as a strategy to reduce data dependency [31, 73, 30, 7, 52]. Instruction tuning refers to fine-tuning a pre-trained model on a curated dataset of instruction-response pairs, which allows the model to follow human directives. While recent efforts have successfully generated large instruction datasets through models like GPT-4V [73], the improvements in Video-LLMs achieved through training on these datasets have been limited. These models often show persistent failure modes in understanding and reasoning about videos, including spatial misalignment [9], temporal incoherence [9], and cross-frame disconnections  [20, 24, 35, 15]. Addressing these alignment challenges through human annotation is prohibitively expensive, as it requires annotators to carefully identify and label examples that demonstrate proper grounding, temporal coherence, and factual consistency. This suggests that merely scaling up models and data may not be the most effective approach. Instead, the key challenge lies in ensuring faithful alignment between model responses and video content.
Figure 1: With just 7k preference pairs, VideoPASTA surpasses the Qwen2.5-VL [3] baseline, LLaVA-Hound [70] and TPO [28] on MVBench, demonstrating that targeted adversarial alignment can outperform larger but less focused datasets.
Recent work has explored video-language alignment [1, 70, 28] using Direct Preference Optimization (DPO) [39]. However, current approaches focus on collecting more preference data rather than addressing the core weaknesses of Video-LLMs, generating data that reinforce existing strengths instead of challenging limitations. Furthermore, these methods rely on proprietary models for generating samples [70] or require video captioning [28], limiting their scalability.
@@
This raises an important research question: How can we align Video-LLMs to understand spatial, temporal, and cross-frame relationships without relying on human annotations, captions, or proprietary models while ensuring computational efficiency? To address this challenge, we introduce VideoPASTA  [Uncaptioned image] (Preference Alignment with Spatio-Temporal-Cross Frame Adversaries), a novel framework that enforces accurate video-language alignment through targeted preference pairs. Our approach contrasts preferred responses with carefully generated adversarial responses that capture three common failure modes in video understanding: (1) spatial misalignment, where responses misrepresent object relationships and interactions within local frames, (2) temporal incoherence, where responses violate the natural progression of events, and (3) cross-frame disconnection, where responses incorrectly link information across distant frames. VideoPASTA leverages DPO and structured preference pairs to address failure modes in video understanding. In summary, our contributions are as follows:
Figure 2: Overview of VideoPASTA Refer to caption. For each aligned query-response pair, we generate three types of targeted adversarial examples: (1) Spatial Mislignment that deliberately confuse object positions and relations (e.g., misplacing the plants relative to the person), (2) Temporal Incoherence that violate event sequences (e.g., claiming simultaneous actions that occur sequentially), and (3) Cross-Frame Disconnection that make invalid connections across distant frames (e.g., incorrectly describing location changes). These pairs are filtered using Qwen2.5-32B [61] to ensure quality, then used to train the model through DPO, maximizing the likelihood gap between preferred and adversarial responses. Refer to caption → trainable, Refer to caption → frozen.
@@
    We present VideoPASTA, a novel DPO-based framework for training video-language models, which addresses three key issues: spatial misalignment, temporal incoherence, and cross-frame disconnection while eliminating the need for human annotation, caption data, or proprietary models.
@@
    We set a new efficiency standard for preference optimization in video-language models, achieving substantial improvements with just 7,020 preference pairs compared to much larger instruction tuning datasets (1.3M) and preference datasets (17k) used in prior work.
@@
    We demonstrate through extensive evaluation across eight video understanding benchmarks that our targeted alignment approach leads to significant relative improvements, including +3.05% on VideoMME, +1.97% on NeXTQA, +1.69% on MVBench, and +1.31% on LongVideoBench compared to baselines.
@@ 2.1 Video-LLMs
Recent advancements have led to the development of numerous Video-LLMs with impressive capabilities [25, 73, 52, 7, 69, 26, 63, 51, 55, 44, 42, 68, 29]. However, comprehensive evaluations [13, 76, 33] reveal persistent challenges in three critical areas. First, temporal reasoning remains a significant hurdle, particularly for long videos. Researchers have attempted to address this through increased context length [32, 42, 69, 43], compression techniques [29, 53, 23, 22, 34], and training-free approaches [62, 18, 59]. While these approaches improve token efficiency, they often fall short in enhancing fundamental temporal understanding. More specialized methods [6, 41, 50, 16, 57] directly target temporal reasoning but require substantial computational resources. Second, spatial misalignment remains a frequent challenge, as models struggle with object localization and spatial relationships, leading to incorrect positioning and poor occlusion handling [40, 4]. Third, cross-frame disconnection often results in failures in maintaining object continuity and narrative coherence across video segments [48, 19]. Most existing methods primarily address one of these dimensions or rely heavily on large-scale instruction-tuning datasets, which fail to resolve the core alignment challenges [52, 51, 73, 30, 5, 71].
VideoPASTA addresses all three failure modes through DPO-based training on structured preference pairs. Rather than addressing isolated weaknesses, we generate preference pairs that challenge the model’s understanding across temporal, spatial, and cross-frame dimensions. This approach enables VideoPASTA to achieve more comprehensive video-language alignment compared to methods that rely on conventional instruction tuning.
2.2 Reward Modeling for Video-LLMs
Reward modeling is essential for aligning VLMs with human preferences by grounding their responses in an accurate visual context. Recent work has proposed a factually augmented RLHF approach that integrates image captions and verified data into the reward function to reduce hallucinations and improve output reliability [47]. Other approaches exploit AI-generated feedback via reinforcement learning to guide video–text alignment, thereby reducing reliance on expensive human annotations [2]. Complementary strategies in reward modeling for VLMs include the conditional preference optimization framework in MDPO [49], the calibrated self-rewarding approach that integrates visual constraints [77], and fine-grained reward modeling via dense sentence-level feedback in ViGoR [60]. In contrast, we employ DPO [39] to directly optimize preference learning, an approach adopted by recent Video-LLMs such as LLaVA-Hound-DPO [70] and i-SRT [1], which relies on proprietary models to generate samples and requires 17k preference pairs.
While previous approaches focus primarily on accumulating diverse preference samples, VideoPASTA fundamentally reimagines preference data generation through systematic targeting of failure modes. By generating preference pairs that deliberately challenge failure modes, we shift towards a preference quality over quantity approach. This reduces the required preference pairs and creates a more robust learning signal that simultaneously improves video understanding across all three critical dimensions.
2.3 Self-Training
To address the challenge of large, annotated datasets, self-training methods [46, 45, 12, 78, 21] have been proposed to use unlabeled or weakly annotated data to improve model performance. However, current methods frequently depend on expensive GPT APIs [45] or require ground truth feedback [78], which limits scalability and hinders comprehensive alignment. While LLaVA-Hound-DPO [70] introduced preference optimization for video understanding, it operates primarily at the text level without directly targeting visual alignment failures and requires 17k preference pairs generated using proprietary models. Similarly, Temporal Preference Optimization (TPO) [28] improves temporal grounding by contrasting predictions from full videos with those from subsampled segments but addresses only temporal aspects while neglecting spatial relationships and cross-frame reasoning. Moreover, TPO still requires video captioning as an intermediate step and uses up to 10k preference pairs, introducing additional computational overhead.
What is missing from current approaches is a unified framework that efficiently targets all three critical dimensions of video understanding without requiring proprietary models, large preference datasets, or intermediate captioning steps. VideoPASTA fills this gap by employing preference pairs that challenge and refine a model’s understanding, achieving superior alignment with just 7k carefully designed preference pairs.
@@ 3 VideoPASTA [Uncaptioned image]
VideoPASTA is a DPO-based framework that helps video-language models stay closely aligned with video content using structured preference optimization. Formally, our approach leverages a preference dataset $D$ defined as follows:
$D=\{(V,q,r^{+},r^{-})\},$ (1)
where $V$ represents the input video, $q$ denotes the query regarding the video content, $r^{+}$ is the preferred response accurately aligned with the video, and $r^{-}$ is the adversarial response deliberately generated to introduce misalignment. Our approach addresses three key failure modes: spatial, temporal, and cross-frame misalignments by generating targeted adversarial examples for each. Specifically, we create one accurate (“aligned”) example that correctly describes a scene, along with three misleading (“adversarial”) examples, each designed to break a different aspect of alignment. By training on these contrastive pairs using DPO, the model learns to better distinguish between preferred responses and adversarial responses. As shown in Figure 2, this enables VideoPASTA to achieve more robust and comprehensive alignment across all aspects of video understanding. We focus on the following three failure modes of video understanding:
(1) Spatial Misalignment.
Models often struggle with spatial relationships over short frame intervals. For example, a model may incorrectly describe a pedestrian as “crossing in front of” a moving vehicle when they are actually walking behind it, a critical spatial error that only becomes apparent through temporal context. Achieving accurate alignment requires understanding object interactions within local temporal contexts.
(2) Temporal Incoherence.
Video frames capture dynamic events. In a cooking demonstration, a model might mistakenly describe a chef dicing vegetables, heating oil, and adding ingredients in a different order. To avoid temporal confusion, models must preserve the correct order of actions and their causal relationships.
(3) Cross-frame Disconnection.
Distant frames provide context for continuity or setting changes. However, models often fail to link events, such as a character entering a building and later appearing in an office, treating them as unrelated scenes instead of a coherent progression. This results in incorrect generalizations and overlooked state changes, underscoring the need for robust cross-frame tracking.
To address each dimension, we generate adversarial responses that intentionally violate spatial, temporal, or cross-frame constraints. By contrasting these adversarial responses with preferred responses using DPO, the model is guided to learn precise alignment for each aspect of video-language understanding.
@@ 3.1 Spatial Misalignment
Spatial relationships, such as object positions, occlusions, depth order, and relative arrangements, are crucial for video understanding. To address these factors, we create targeted preference pairs that focus on spatial alignment.
@@
We use InternVL2.5-38B [7] to automatically generate a variety of spatial queries (prompt provided in Appendix, Figure 10). These queries cover key aspects of spatial understanding, including occlusion (e.g., “Which object is partially hidden behind another?”), depth perception (e.g., “Which item appears closest to the camera?”), relative positioning (e.g., “How many objects occupy the left vs. right third of the frame?”), foreground-background relationships, and frame layout (e.g., “Which objects are at the top versus bottom edge?”).
Preferred Response Generation.
To capture fine-grained spatial details, we sample the input video at 32 fps. The baseline Qwen2.5-VL [3] model is then explicitly prompted to provide coherent and detailed descriptions of object interactions, depth cues, and spatial configurations across neighboring frames. This ensures that the generated preferred responses accurately reflect the true spatial relationships in the video.
Adversarial Response Generation.
We generate challenging examples by undersampling videos at 1 fps and modifying prompts to induce spatial errors. For occlusion queries, adversarial prompts instruct the baseline model to describe all objects as fully visible, even when occluded (e.g., stating “the person is fully visible” when partially hidden behind a counter). For depth-related queries, prompts require the model to reason that all objects are equidistant, disregarding clear perspective cues (e.g., claiming “the cup and the mountain are at the same distance from the camera”). These intentionally flawed responses act as adversaries to test and refine precise spatial reasoning.
@@ 3.2 Temporal Incoherence
Accurately capturing event sequences, action transitions, and causal relationships is crucial for robust video understanding. To evaluate and enhance temporal reasoning, we create preference pairs that focus on the model’s temporal coherence.
@@
We use InternVL2.5-38B to generate queries that focus on key aspects of temporal understanding (prompt provided in Appendix, Figure 11). These include event ordering (e.g.,“Which major action occurs first, and which follows?”), action boundaries (e.g., “Does the person complete one task before starting the next?”), transition points (e.g., “When does the subject switch activities?”), causality (e.g., “Is the second event a direct result of the first?”), and concurrent actions (e.g., “Are there any simultaneous events, and how do they overlap?”).
Preferred Response Generation.
To obtain temporally coherent responses, we process the video at its native frame rate, ensuring dense temporal sampling. The baseline model is prompted to describe the precise sequence of events, effectively capturing transitions, dependencies, and causal relationships.
Adversarial Response Generation.
We create adversarial examples by undersampling the videos at 1 fps and utilizing prompts that induce temporal distortions. One adversarial prompt instructs the baseline model to describe sequential actions as occurring simultaneously (e.g., “the chef chopping vegetables, heating the pan, and plating the dish all at the same time”). Another prompts the model to ignore action transitions, merging distinct events into a continuous sequence (e.g., describing “a continuous swimming motion” when the video shows separate diving, swimming, and exiting phases). These flawed responses act as adversaries to test and refine precise temporal reasoning.
@@ 3.3 Cross-frame Disconnection
Robust video understanding requires capturing long-range cross-frame relationships, such as object continuity, character persistence, setting changes, and narrative progression across distant frames. To achieve this, we create targeted preference pairs that focus specifically on these cross-frame dependencies.
@@
We use InternVL2.5-38B to automatically generate queries that evaluate various aspects of cross-frame understanding (prompt provided in Appendix, Figure 12). For instance, queries ask whether the same object appears in both the opening and closing scenes, whether characters present early in the video reappear, whether the setting evolves over time, whether actions repeat across distant segments, and how early events foreshadow later developments.
Preferred Response Generation.
To capture coherent long-range dependencies, we uniformly sample the entire video sequence at its native frame rate. The baseline model is prompted to describe consistent object transformations, character developments, setting changes, and narrative connections, ensuring that preferred responses accurately reflect the continuity and narrative progression throughout the video.
Adversarial Response Generation.
We generate challenging examples by undersampling the video at 1 fps by using prompts that disrupt cross-frame connections. For object continuity, adversarial prompts instruct the baseline model to treat identical objects in different scenes as unrelated (e.g., describing “a new red car appears” when it is the same vehicle shown from a different angle). For character persistence, prompts require the model to treat similar-looking characters as entirely distinct (e.g., stating “a different person enters the office” when it is clearly the same individual from earlier). These flawed responses act as adversaries to test and strengthen robust, long-range video-language alignment.

### v3

@@ Abstract
Video-language models (Video-LLMs) excel at understanding video content but struggle with spatial relationships, temporal ordering, and cross-frame continuity. To address these limitations, we introduce VideoPASTA  [Uncaptioned image] (Preference Alignment with Spatio-Temporal-Cross Frame Adversaries), a framework that enhances Video-LLMs through targeted preference optimization. VideoPASTA trains models to distinguish accurate video representations from carefully crafted adversarial examples that deliberately violate spatial, temporal, or cross-frame relationships. With only 7,020 preference pairs and Direct Preference Optimization, VideoPASTA enables models to learn robust representations that capture fine-grained spatial details and long-range temporal dynamics. Experiments demonstrate that VideoPASTA is model agnostic and significantly improves performance, for example, achieving gains of up to +3.8 percentage points on LongVideoBench, +4.1 on VideoMME, and +4.0 on MVBench, when applied to various state-of-the-art Video-LLMs. These results demonstrate that targeted alignment, rather than massive pretraining or architectural modifications, effectively addresses core video-language challenges. Notably, VideoPASTA achieves these improvements without any human annotation or captioning, relying solely on 32-frame sampling. This efficiency makes our approach a scalable plug-and-play solution that seamlessly integrates with existing models while preserving their original capabilities.
VideoPASTA [Uncaptioned image]: 7K Preference Pairs That Matter for Video-LLM Alignment
Yogesh Kulkarni    Pooyan Fazli Arizona State University https://people-robots.github.io/VideoPASTA/
Figure 1: With just 7k preference pairs, VideoPASTA outperforms the Qwen2.5-VL Bai et al. (2025) baseline, LLaVA-Hound Zhang et al. (2025), and TPO Li et al. (2025b) on MVBench, showing that targeted alignment surpasses models trained on larger datasets.
@@ 1 Introduction
Recent advances in video language models (Video-LLMs) have enabled efficient video understanding and reasoning, achieving strong performance on tasks like captioning and question answering Li et al. (2025a); Wang et al. (2022, 2024d); Bai et al. (2025); Chen et al. (2024c). However, these models typically rely on large, high-quality annotated datasets and significant computing resources for training. Instruction tuning has emerged as a way to reduce data requirements by fine-tuning models on curated instruction-response pairs Liu et al. (2023); Zhang et al. (2024d); Lin et al. (2024); Chen et al. (2024c); Wang et al. (2024b). While large instruction datasets have been generated using models like GPT-4V Zhang et al. (2024d), improvements from training on these datasets remain limited. Video-LLMs still struggle with spatial misalignment, temporal incoherence, and cross-frame disconnections Choong et al. (2024); Huang et al. (2025); Leng et al. (2024); Ma et al. (2024); Gunjal et al. (2024). Addressing these issues through human annotation is expensive, as it requires identifying examples with proper grounding and coherence. This suggests that merely scaling models and data alone is insufficient. Instead, the core challenge lies in achieving faithful alignment between model outputs and video content. Recent work has applied Direct Preference Optimization (DPO) to improve video-language alignment Ahn et al. (2024a); Zhang et al. (2025); Li et al. (2025b); Rafailov et al. (2023). However, these methods often reinforce existing strengths through collecting more preference data instead of targeting core weaknesses in Video-LLMs. They also rely on proprietary models Zhang et al. (2025) or video captioning Li et al. (2025b), limiting scalability.
@@
This raises a key question: How can we align Video-LLMs to understand spatial, temporal, and cross-frame relationships without human annotations, captions, or proprietary models, while remaining computationally efficient? To address this, we introduce VideoPASTA  [Uncaptioned image] (Preference Alignment with Spatio-Temporal-Cross Frame Adversaries), a model-agnostic framework that improves video-language alignment using targeted preference pairs. VideoPASTA contrasts aligned (“preferred”) responses with adversarial ones that capture three common failure modes in video understanding: (1) spatial misalignment, where responses misrepresent object relationships and interactions, (2) temporal incoherence, where responses violate the natural progression of events, and (3) cross-frame disconnection, where responses violate object persistence, character consistency, and narrative progression across more distant parts of a video. By combining DPO with structured preference data, VideoPASTA directly tackles these limitations in Video-LLMs. In summary, our contributions are as follows:
@@
    We introduce VideoPASTA, a novel, model-agnostic DPO framework that improves video-language alignment by addressing spatial misalignment, temporal incoherence, and cross-frame disconnection, without human annotations, captions, or proprietary models.
@@
    VideoPASTA sets a new efficiency benchmark, achieving strong results using just 7,020 preference pairs, far fewer than prior instruction tuning (1.3M) or preference datasets (17k).
@@
    Extensive evaluations on seven benchmarks show consistent, model-agnostic gains, with improvements of up to +3.8 percentage points on LongVideoBench, +4.1 on VideoMME, and +4.0 on MVBench.
Figure 2: Overview of VideoPASTA Refer to caption. For each aligned query, we generate three types of targeted adversarial examples: (1) Spatial Misalignment, which intentionally distorts object positions or relationships (e.g., misplacing the plants relative to the person); (2) Temporal Incoherence, which violates event order (e.g., describing sequential actions as occurring simultaneously); and (3) Cross-Frame Disconnection, which introduces incorrect links across distant frames (e.g., misrepresenting location changes). We filter these pairs using Qwen2.5-32B Yang et al. (2024) to ensure quality and use them to train the model via DPO, optimizing for a larger likelihood gap between aligned and adversarial responses. Refer to caption → trainable, Refer to caption → frozen.
@@ 2 Related Work
Despite advances in Video-LLMs Li et al. (2024a); Wang et al. (2024b); Chen et al. (2024c), evaluations Fu et al. (2025); Zhou et al. (2025); Liu et al. (2024b) reveal persistent challenges in three key areas. First, temporal reasoning, especially in long videos, remains difficult. Approaches like longer context Shen et al. (2024); Zhang et al. (2024b), compression Li et al. (2024c); Wang et al. (2024c), and training-free methods Yang et al. (2025); Huang et al. (2025) improve token efficiency but not core understanding, while specialized methods Chen et al. (2024b); Ren et al. (2024) demand heavy computation. Second, spatial misalignment leads to poor object localization and occlusion handling Ranasinghe et al. (2024); Chen et al. (2024a). Third, cross-frame disconnection disrupts continuity and narrative coherence Tan et al. (2024); Huang et al. (2024). Most methods address only one issue or rely on large-scale instruction tuning Wang et al. (2024b); Zhang et al. (2024d); Lin et al. (2024), which fails to solve these core alignment problems. VideoPASTA uses DPO-based training on structured preference pairs to jointly address temporal, spatial, and cross-frame failures. By challenging models across all three dimensions, it achieves more comprehensive video-language alignment than conventional instruction tuning.
Video-Language Alignment.
While reward modeling (Sun et al., 2024; Ahn et al., 2024b; Wang et al., 2024a) and self-training methods (Deng et al., 2024; Zohar et al., 2025; Kulkarni and Fazli, 2025b, a) aim to improve video-language alignment and reduce the need for manual annotations, existing approaches still face major limitations. Prior DPO applications, such as LLaVA-Hound-DPO (Zhang et al., 2025) and i-SRT (Ahn et al., 2024a), often depend on proprietary models to generate training data, require large-scale preference datasets (e.g., 17k pairs), and focus mainly on text-level alignment rather than visual grounding. Other methods, like Temporal Preference Optimization (TPO) (Li et al., 2025b), target only one dimension, such as temporal reasoning, using up to 10k pairs and relying on intermediate captioning, while overlooking spatial and cross-frame aspects. This highlights the need for a unified framework that efficiently addresses all three key failure modes without relying on costly dependencies. VideoPASTA addresses this gap by generating just 7k carefully designed preference pairs that explicitly challenge a model’s spatial, temporal, and cross-frame understanding. This “quality over quantity” approach avoids the need for human annotations, captions, or proprietary models, delivering a stronger and more efficient learning signal for comprehensive Video-LLM alignment.
@@ 3 VideoPASTA [Uncaptioned image]
VideoPASTA is a DPO-based framework designed to improve Video-LLM alignment by addressing three key failure modes: spatial misalignment, temporal incoherence, and cross-frame disconnection. It optimizes over a structured preference dataset $D=\{(V,q,r^{+},r^{-})\}$, where $V$ is the input video, $q$ is a query, $r^{+}$ is the aligned (preferred) response, and $r^{-}$ is an adversarial (misleading) response that introduces deliberate misalignment.
For each input video, we generate an aligned query $q$. We then produce a aligned response $r^{+}$ using densely sampled frames (32fps) and an adversarial response $r^{-}$ using sparsely sampled frames (1fps), elicited by a deliberately flawed adversarial query. This adversarial query is only a data generation tool; the final training triplet remains $(V,q,r^{+},r^{-})$, pairing the adversarial response with the original aligned query. The 32:1 sampling ratio is a design choice to ensure factual accuracy in aligned responses while inducing errors in adversarial ones. Training on these pairs via DPO improves the model’s video-language alignment. Figure 2 shows the full pipeline. All prompts are provided in the Appendix (Figures 91011, and  12).
@@ 3.1 Spatial Misalignment
We create targeted preference pairs that focus on spatial alignment.
@@
We generate a variety of spatial queries covering key aspects of spatial understanding, including occlusion (e.g., “Which object is partially hidden?”), depth perception (e.g., “Which item appears closest to the camera?”), relative positioning (e.g., “How many objects are present in the left third vs. the right third of the frame?”), foreground-background relationships, and frame layout (e.g., “Which objects are near the top edge vs. the bottom edge?”).
Response Generation.
We generate aligned responses from videos sampled at their native frame rate to capture fine-grained spatial details. This ensures that the generated aligned responses accurately reflect the true spatial relationships in the video. Corresponding adversarial responses are generated using prompts specifically designed to induce spatial errors (e.g., describing occluded objects as fully visible or ignoring depth cues).
@@ 3.2 Temporal Incoherence
To enhance temporal reasoning, we create preference pairs that focus on the model’s temporal coherence.
@@
We generate queries focused on key temporal aspects, including event ordering (e.g., “What occurs first?”), action boundaries (e.g., “Does the person complete one task before starting the next?”), transition points (e.g., “When does the subject switch activities?”), and causality (e.g., “Is the second event a direct result of the first?”).
Response Generation.
We generate aligned responses that accurately describe the sequence of events, effectively capturing transitions, dependencies, and causal relationships. Corresponding adversarial responses are generated using prompts that induce temporal distortions (e.g., describing sequential actions as simultaneous or merging distinct events).
@@ 3.3 Cross-Frame Disconnection
Robust video understanding requires capturing long-range cross-frame relationships. We create targeted preference pairs that focus on these dependencies.
@@
We generate queries focused on cross-frame dependencies, including object continuity (e.g., “Does the same object reappear in both the opening and closing scenes?”) and narrative links (e.g., “Do early events foreshadow later developments?”).
Response Generation.
We generate aligned responses that accurately reflect object transformations, character continuity, setting changes, and narrative flow. Adversarial responses are generated using prompts that intentionally break these continuities (e.g., describing “a new red car appears” when it is the same vehicle from a different angle).